OpenStudy (anonymous):

Solve the following system of equations graphically. Click on the graph until the correct one appears. 2x -y = -1 x + y = -2

OpenStudy (callisto):

2x -y = -1 -(1) x + y = -2 -(2) (1) + (2) 3x = -3 x = ... Put x = ... into (2) y = -2 - x = ???

OpenStudy (callisto):

Then, match the point with the graph and you'll get your answer!~

OpenStudy (anonymous):

ok